For each of the following accounts, give the growth factor per compounding period, then give the annual growth factor and the annual percent change (APY).
Account A has a 5% APR compounded monthly.
Monthly growth factor: Correct
Annual growth factor: Correct
APY: Incorrect %
Account B has a 3.2% APR compounded daily. Assume there are 365 days in the year.
Daily growth factor: Incorrect
Annual growth factor: Incorrect
APY: Incorrect %

Answer by greenestamps(13200)



It doesn't make sense to post a question like this and only tell us this answer is correct and this other answer is incorrect, without showing those answers.

If you want help in understanding these problems, SHOW US your correct and incorrect answers so we know HOW we can help.

The basics are quite simple if you know the vocabulary.

Example: 6% compounded quarterly (4 times per year).

APR: 6% = .06
quarterly interest rate: .06/4 = .015
quarterly growth factor: 1+.015 = 1.015
yearly growth factor: (1.015)^4 = 1.06136 to 5 decimal places
APY: 1.06136-1 = .06136 = 6.136%